# Haar-Type Measures on Topological Quasigroups and Kunen’s Theorem

### 06 Mar 2026 v1

@@ 3 The Moufang Identity and Translation Calculus
Theorem 1 (Translation form of the Moufang identity).
@@ 10 Existence Questions and the Role of the Translation Group
11 Toward a Measure-Theoretic Reading of Kunen’s Theorem
We can now formulate the main conceptual point of the paper.
Conjecture 1.
Let $Q$ be a locally compact Hausdorff topological quasigroup satisfying $(N1)$. Suppose that $Q$ admits a nonzero two-sided quasi-invariant measure with modular cocycles $j$ and $\rho$. Then the identity $(N1)$ forces the left cocycle $j$ to satisfy strong rigidity conditions, and under suitable additional assumptions this rigidity collapses $j$ to the trivial cocycle:
$j(a)=1\qquad(a\in Q).$
If such a collapse can be established, then the quasigroup would become, from the measure-theoretic point of view, analogous to a unimodular object. The guiding philosophy is that this unimodularity should correspond to the emergence of loop structure.
@@ 11 Toward a Measure-Theoretic Reading of Kunen’s Theorem
At present this is a research program rather than a completed proof of Kunen’s theorem. The point of the present note is to make precise the modular-cocycle framework and to show that the identity $(N1)$ interacts with it in a nontrivial and structurally meaningful way.
@@ 12 Translation–cocycle compatibility
By Theorem LABEL:thm:translation-form-N1, the Moufang identity $(N1)$ is equivalent to the operator identity
@@ 14 Unsolved Problems
15 Existence Issues and the Unimodularity Perspective
@@ 15.3 Unimodularity and loop structure
Conjecture 2.
Let $Q$ be a locally compact topological quasigroup satisfying the Moufang identity. If a quasi-invariant measure exists and the associated modular cocycle satisfies suitable regularity conditions, then the cocycle collapses to the trivial one
$j\equiv 1.$
Such a result would suggest that loop structure may be interpreted as a form of unimodularity in the translation geometry of quasigroups.
@@ 16 Conclusion
We introduced a framework for Haar-type measures on topological quasigroups, centered on quasi-invariance under translations and the associated modular cocycles. We then showed, under explicit two-sided quasi-invariance assumptions, that the Moufang-type identity $(N1)$ yields a rigorous multiplicativity relation for the left cocycle.
This places the quasigroup situation in close analogy with the classical theory of the modular function on locally compact groups. In particular, it supports the idea that loop structure may be interpreted as a kind of unimodularity in the translation geometry of a quasigroup.
Several problems remain open, most importantly the existence theory for quasi-invariant measures on topological quasigroups and the precise mechanism by which cocycle rigidity should force an identity element. These questions suggest that the interaction among quasigroup theory, topology, and harmonic analysis deserves further study.
@@
Appendix A Structural Outline of the Theory
For the convenience of the reader we summarize here the logical structure of the theory developed in this paper. The framework connects quasigroup structure, translation operators, and measure-theoretic cocycles.
Algebraic foundation
  • Definition A.1 (Quasigroup). A quasigroup is a set $Q$ with a binary operation such that for every $a,b\in Q$ the equations
    $ax=b,\qquad ya=b$
    have unique solutions.
  • Definition A.2 (Topological quasigroup). A topological quasigroup is a quasigroup endowed with a Hausdorff topology such that multiplication and division operations are continuous.
Translation operators
  • Definition A.3 (Translations).
    $L_{a}(x)=ax,\qquad R_{a}(x)=xa.$
  • Definition A.4 (Multiplication groups).
    $\mathrm{LMlt}(Q)=\langle L_{a}\mid a\in Q\rangle,\qquad\mathrm{Mlt}(Q)=\langle L_{a},R_{a}\mid a\in Q\rangle.$
Measure-theoretic framework
  • Definition A.5 (Pushforward measure).
    For a measurable map $T:X\to X$ and a measure $\mu$,
    $(T_{*}\mu)(E)=\mu(T^{-1}(E)).$
  • Lemma A.6 (Functoriality of pushforward).
    $(S\circ T)_{*}\mu=S_{*}(T_{*}\mu).$
  • Definition A.7 (Quasi-invariant measure).
    A measure $\mu$ on $Q$ is left quasi-invariant if
    $(L_{a})_{*}\mu=j(a)\mu,$
    and right quasi-invariant if
    $(R_{a})_{*}\mu=\rho(a)\mu.$
Translation form of Moufang identity
  • Theorem A.8 (Translation form of the Moufang identity).
    The Moufang identity
    $((xy)z)y=x(y(zy))$
    is equivalent to
    $R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}.$
Modular cocycles
  • Theorem A.9 (Cocycle multiplicativity).
    Let $\mu$ be a two-sided quasi-invariant measure satisfying
    $(L_{a})_{*}\mu=j(a)\mu,\qquad(R_{a})_{*}\mu=\rho(a)\mu.$
    If $Q$ satisfies the Moufang identity, then the modular cocycle satisfies the compatibility relation
    $j(xy)\rho(y)=j(x)j(y)\rho(y).$
    In particular, since $\rho(y)>0$, we obtain
    $j(xy)=j(x)j(y).$
Multiplicative functions
  • Definition A.10 (Multiplicative function).
    A function
    $\chi:Q\to\mathbb{R}_{>0}$
    is multiplicative if
    $\chi(xy)=\chi(x)\chi(y).$
  • Lemma A.11 (Normalization).
    If $Q$ has an identity element $e$, then
    $\chi(e)=1.$
Representation of the translation group
  • Theorem A.12 (Translation representation).
    If $\chi:Q\to\mathbb{R}_{>0}$ is multiplicative, then
    $L_{a}\mapsto\chi(a)$
    extends to a homomorphism
    $\pi:\mathrm{LMlt}(Q)\to\mathbb{R}_{>0}.$
Conceptual structure
The logical structure of the theory can be summarized schematically as
quasigroup structure
$\downarrow$
translation operators
$\downarrow$
quasi-invariant measures
$\downarrow$
modular cocycles
$\downarrow$
Moufang identity
$\downarrow$
cocycle multiplicativity
$\downarrow$
multiplicative functions
$\downarrow$
$\text{representation of the translation group}.$
This framework suggests a measure-theoretic viewpoint in which loop structure may be interpreted as a form of unimodularity in the translation geometry of quasigroups.
Appendix B Structural Diagram of the Theory
For the convenience of the reader, we summarize here the logical architecture of the paper in diagrammatic form.
@@ Appendix B Structural Diagram of the Theory
The guiding interpretation is that the modular cocycle measures the defect of translation invariance, while the Moufang identity imposes a rigidity condition on this defect. In this sense, loop structure may be viewed as a form of unimodularity in the translation geometry of quasigroups.

### 11 Mar 2026 v2

@@ Abstract
This perspective suggests a possible measure-theoretic interpretation of Kunen’s theorem in terms of unimodularity of the translation geometry.
@@ 3 The Moufang Identity and Translation Calculus
Theorem 1 (Translation form of the Moufang identity (N1)).
@@ 10 Existence Questions and the Role of the Translation Group
11 Towards a Measure-Theoretic Interpretation of Kunen’s Theorem
The purpose of this paper is to develop a framework in which Haar-type structures on topological quasigroups can be studied via quasi-invariant measures and modular cocycles. From this measure-theoretic viewpoint, loop identities may be interpreted as a form of unimodularity in the translation geometry of quasigroups.
This perspective naturally leads to the following conjecture.
Conjecture 1 (Modular Collapse Conjecture).
Let $Q$ be a locally compact Hausdorff topological quasigroup satisfying the Moufang-type identity $(N1)$. Assume that $Q$ admits a quasi-invariant measure with modular cocycle $j$.
Then the cocycle collapses:
$j\equiv 1.$
In particular, the translation geometry becomes unimodular in the measure-theoretic sense.
This conjecture is consistent with Kunen’s theorem, which shows that the identity $(N1)$ already imposes strong structural constraints on quasigroups.
@@ 11 Towards a Measure-Theoretic Interpretation of Kunen’s Theorem
A similar conjecture should hold under a Moufang identity in place of $(N1)$. The identity $(N1)$ is emphasized here because of its connection with Kunen’s theorem.
This conjecture suggests that Kunen’s theorem may admit a measure-theoretic interpretation in terms of unimodularity of the translation geometry.
@@ 12 Translation–cocycle compatibility
By Theorem 1, the Moufang identity $(N1)$ is equivalent to the operator identity
@@ 14 Unsolved Problems
15 Existence Issues and
the Unimodularity Perspective
@@ 16 Conclusion
We introduced a framework for studying Haar-type structures on topological quasigroups using quasi-invariant measures and modular cocycles. Under explicit two-sided quasi-invariance assumptions, we showed that the Moufang-type identity $(N1)$ yields a rigorous multiplicativity relation for the left cocycle.
This places the quasigroup situation in close analogy with the classical theory of the modular function on locally compact groups.
Several problems remain open, most importantly the existence theory for quasi-invariant measures on topological quasigroups and the precise mechanism by which cocycle rigidity might force the existence of an identity element. These questions suggest that the interaction among quasigroup theory, topology, and harmonic analysis deserves further study.
@@
Remark 15.
In this paper we proposed a framework for studying Haar-type structures on topological quasigroups using quasi-invariant measures and modular cocycles. This perspective suggests that loop identities may admit a measure-theoretic interpretation in terms of unimodularity of the translation geometry. In particular, the conjectural collapse of the modular cocycle under the identity $(N1)$ would provide a new viewpoint on Kunen’s theorem from the perspective of harmonic analysis on quasigroup-like structures.
Future directions.
Several natural questions arise from the present framework. First, it would be important to develop a general existence theory for quasi-invariant measures on locally compact topological quasigroups. Second, one may ask under what structural conditions the rigidity of the modular cocycle forces the existence of an identity element, thereby producing loop structure. These problems suggest that the interaction between quasigroup theory, topology, and harmonic analysis deserves further investigation.
Appendix A Structural Diagram of the Measure-Theoretic Framework
For the convenience of the reader, we summarize the conceptual architecture of the paper.
Interpretation.
The diagram clarifies the conceptual mechanism developed in this paper. The Moufang-type identity $(N1)$ constrains the translation calculus of the quasigroup. When a quasi-invariant measure exists, this constraint propagates to the modular cocycle and forces a rigidity relation. The resulting multiplicativity of the cocycle suggests a representation-theoretic interpretation of the translation geometry. In this sense, the emergence of loop structure in Kunen’s theorem may be interpreted as a collapse of the modular defect.
@@ Appendix A Structural Diagram of the Measure-Theoretic Framework
Kunen collapse mechanism.
Kunen’s theorem shows that the Moufang-type identity $(N1)$ imposes strong constraints on the translation structure of a quasigroup. These constraints force the existence of a two-sided identity element, and hence the quasigroup collapses to a loop. The diagram illustrates this structural mechanism.
Quasigroup QQMoufang-type identity (N1)(N1)Translation constraintsExistence of a two-sided identityLoop structure